# GNSS-based Lunar Orbit and Clock Estimation With Stochastic Cloning UD Filter

### v2

@@ 7.2.5 Estimation Error Distribution
These trends are also reflected in the SISE histograms over the final orbit across 50 Monte-Carlo runs (Figs. 16 and 17). Incorporating TDCP shifts both the position and velocity SISE distributions toward smaller values for both filtering and smoothing. Comparison of filter versus smoother results shows that smoothing reduces the tails of the SISE distributions, consistent with its ability to incorporate future measurements to refine poorly observed portions of the orbit.
@@ 8 Conclusion
We evaluated the proposed approach using high-fidelity Monte-Carlo simulations that incorporate realistic multi-constellation GNSS geometry, broadcast-ephemeris errors, and ray-tracing-based ionospheric/plasmaspheric delay models. The results show that augmenting ionosphere-free pseudorange with TDCP consistently improves estimation performance relative to pseudorange-only baselines. In particular, TDCP reduces the filtering RMS position and velocity SISE by approximately 16–18% and 9–11%, respectively, achieving $3.68\text{\,}\mathrm{m}$ RMS position SISE and $0.39\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ RMS velocity SISE under the tested configuration. The delayed-state smoother further reduces residual errors and tails in the error distribution, yielding sub-$4\text{\,}\mathrm{m}$ position accuracy and sub-$\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ velocity accuracy, which satisfies the stringent accuracy targets envisioned for future LANS.
Future work will focus on (i) systematic tuning of tangential-altitude masking and measurement noise inflation to better trade measurement availability against plasmaspheric-delay biases, and (ii) reducing reliance on masking by explicitly estimating residual ionospheric/plasmaspheric terms and multi-constellation inter-system biases within the state (or as consider states). We will also investigate processing raw pseudorange and GRAPHIC observables with ambiguity handling to improve robustness and availability for single-frequency or intermittently dual-frequency reception scenarios.
@@ Acknowledgments
This work was supported in part by funding from the the Nakajima Foundation. The authors would like to thank Kaila Coimbra, Adam Dai, and members of the Stanford NAV Lab for valuable discussions and technical feedback. AI models (ChatGPT, Gemini) were used to assist with language editing and technical clarity of the manuscript. All derivations, figures, simulations, and conclusions were developed, verified, and validated solely by the authors.

### v3

@@ 7.2.5 Estimation Error Distribution
These trends are also reflected in the SISE histograms over the final orbit across 40 Monte-Carlo runs (Figs. 16 and 17). Incorporating TDCP shifts both the position and velocity SISE distributions toward smaller values for both filtering and smoothing. Comparison of filter versus smoother results shows that smoothing reduces the tails of the SISE distributions, consistent with its ability to incorporate future measurements to refine poorly observed portions of the orbit.
@@ 8 Conclusion
We evaluated the proposed approach using high-fidelity Monte-Carlo simulations that incorporate realistic multi-constellation GNSS geometry, broadcast-ephemeris errors, and ray-tracing-based ionospheric and plasmaspheric delay models. The results show that augmenting ionosphere-free pseudorange with TDCP yields a clear improvement in estimation performance relative to pseudorange-only baselines. For filtering, the inclusion of TDCP reduces the RMS position SISE from $6.05\text{\,}\mathrm{m}$ to $4.53\text{\,}\mathrm{m}$, corresponding to an improvement of approximately 25%, while the RMS velocity SISE remains comparable at about $0.45\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$.
The delayed-state smoother provides additional gains by reducing both the mean error levels and the tails of the error distributions. With ionosphere-free pseudorange and TDCP, the smoother achieves an RMS position SISE of $2.79\text{\,}\mathrm{m}$ and an RMS velocity SISE of $0.30\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$, while substantially suppressing high-percentile velocity errors induced by occasional TDCP outliers. These results demonstrate that the combined use of TDCP and smoothing enables sub-$3\text{\,}\mathrm{m}$ position accuracy and sub-$\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ velocity accuracy, meeting the position accuracy target envisioned for LANS.
[…]
@@ Acknowledgments
This work was supported in part by funding from the the Nakajima Foundation. The authors would like to thank Kaila Coimbra, Adam Dai, and members of the Stanford NAV Lab for valuable discussions and technical feedback, and Rebecca Wang and Todd Walter for their advice on GNSS ephemeris error modeling. AI models (ChatGPT, Gemini) were used to assist with language editing and technical clarity of the manuscript. All derivations, figures, simulations, and conclusions were developed, verified, and validated solely by the authors.